### v3

@@ 5.3 NNP MD (NequIP/LAMMPS)
The NequIP plugin for LAMMPS[85] was used to perform several NVT simulations at 300 K for over a nanosecond each. A Nosé-Hoover thermostat was attached to every degree of freedom to ensure equilibration [81]. The long-range Coulomb interactions were added to the simulation using the LAMMPS hybrid overlay method. No initial data was discarded, as the initial frame was taken from the end of the AIMD simulation or classical simulation.
@@
The total simulation times were 8.4 ns for the 2.5 M LiCl, 12 ns for the 1.3 M. 5.3 ns for the 2.5 M KCl and 3.5 ns for the 6.6 M. For LiBr 2.7 ns.
@@ 5.4 Kirkwood-Buff theory calculations
$\frac{1}{1+\rho\left(G_{\text{cc}}-G_{\text{co}}\right)}$ (1)
@@ 5.5 Diffusion coefficients calculation
The results were finally adjusted by finite size corrections.[88] Here, we have used the experimental value (0.888 mPas) for the viscosity of pure water when determining the finite size correction. Experimental values were obtained from Ref. 67.
(a) 4M
(b) 2.6 M
(c) 1.3M
Figure 10: CG model trained on the classical data compared with all-atom classical MD simulations for NaCl.
(a) Lithium water RDFs in LiCl.
(b) Chloride water RDFs in LiCl
(c) Potassium water RDFs.
(d) Bromide water RDFs
Figure 11: Ion water RDFs at 2.5 M
@@ 5.6 Classical force field
The classical molecular dynamic simulation was performed using LAMMPS program, applying the Dang-Smith[89] Lennard-Jones force field. Cross interactions were computed using the Lorentz-Berthelot mixing rule. A 5 ns simulation was conducted with an NVT ensemble at 300 K, controlled by the Nose-Hover thermostat. The simulation was carried out in a box with a volume of 29.83 Å3, using a periodic boundary conditions. The system contained 64 sodium ions, 64 chloride ions and 810 SPC/E water molecules, representing a 4M NaCl solution. Long-range Coulombic interactions were calculated by particle-particle particle-mesh method with a relative force set as 10-5. The cut-off was set 15 Å , slightly extended to improve accuracy in long-range Coulombic interactions.
5.7 Classical coarse-grained model
The classical model was used to provide additional testing of the coarse-grained model. The process is similar to that described for first principle coarse-grained model. Initially, the training data was generated from a 5 ns classical run of a 4M NaCl solution by rerunning the classical trajectory and extracting only ions data every 20 frames, resulting in a total of 25000 frames. Again the long-range electrostatic interactions were removed from forces and energies in the dataset. In NNP training, all the hyperparameters were set as previously mentioned, except for a longer radial cutoff of 15 Å and higher maximum $l$ of 2. Finally, we simulated NaCl at 1.3M, 2.6M and 4M using thbe coarse-grained MD simulations with the force field trained on the classical MD simulation of NaCl at 4M. All other classical coarse-grained settings followed with previous method.
5.8 Ion water RDFs
The full ion water RDFs are shown in Figure 11.
5.9 Lithium dimer forces
To confirm that the lithium dimer formed is physically reasonable we recomputed the forces on the lithium ions in the dimer conformation and compared with the NNP predictions acheiving good agreement (Figure 12).
Figure 12: Forces on the lithium ions in the lithium dimer computed with DFT compared with NNP predictions showing good agreement.

### v4

@@ 5.3 NNP MD (NequIP/LAMMPS)
The NequIP plugin for LAMMPS[85] was used to perform several NVT simulations at 300 K for over a nanosecond each. A Nosé-Hoover thermostat was attached to every degree of freedom to ensure equilibration [81]. The long-range Coulomb interactions were added to the simulation using the LAMMPS hybrid overlay method. No initial data was discarded, as the initial frame was taken from the end of the FPMD simulation or classical simulation.
@@
The total simulation times were 8.4 ns for the 2.5 M LiCl; 12 ns for the 1.3 M LiCl; 5.3 ns for the 2.5 M KCl; 3.5 ns for the 6.6 M KCl; 2.7 ns for LiBr 2.5 M.
@@
We also include a small Gaussian repulsion in the OH interaction to prevent occasional hydrogen dissociation from water. This only acts in the 0 density region so should have no effect on the simulations.
The coarse grained NNPs were also used for molecular dynamics simulations with identical parameters to the all atom MD simulations. These were run for 6 million steps (3 ns).
The KCl slab simulation contained 216 KCl inintialised in a crystal structure, in a 18.62Åx 18.62Å28.62Åcell, where there was a 10 Åvacuum region.
@@ 5.4 Kirkwood-Buff theory calculations
$\frac{d\ln\gamma}{d\ln c}=\frac{1}{1+\rho\left(G_{\text{cc}}-G_{\text{co}}\right)}$ (1)
@@ 5.5 Diffusion coefficients calculation
where the MSD is:
$\text{MSD}(t)=\frac{1}{N}\sum_{i=1}^{N}\left|\mathbf{r}_{i}(t)-\mathbf{r}_{i}(0)\right|^{2}$ (6)
and $t$ is time. The results were finally adjusted by finite size corrections. (See eq. 12 of Ref. 88.) Here, we have used the experimental value (0.888 mPas) for the viscosity of pure water when determining the finite size correction. Experimental values were obtained from Ref. 89.
@@ 5.6 Classical force field
The classical molecular dynamic simulation was performed for NaCl using the LAMMPS program, applying the Dang-Smith[90] Lennard-Jones force field. Cross interactions were computed using the Lorentz-Berthelot mixing rule. A 5 ns simulation was conducted with an NVT ensemble at 300 K, controlled by the Nose-Hover thermostat. The simulation was carried out in a box with a volume of 29.83 Å3, using a periodic boundary conditions. The system contained 64 sodium ions, 64 chloride ions and 810 SPC/E water molecules, representing a 4M NaCl solution. Long-range Coulombic interactions were calculated by particle-particle particle-mesh method with a relative force set as 10-5. The cut-off was set 15 Å , slightly extended to improve accuracy in long-range Coulombic interactions.
The classical model was used to provide additional testing of the coarse-grained model. The process is similar to that described for first principle coarse-grained model. Initially, the training data was generated from a 5 ns classical run of a 4M NaCl solution by rerunning the classical trajectory and extracting only ions data every 20 frames, resulting in a total of 25000 frames. Again the long-range electrostatic interactions were removed from forces and energies in the dataset. In NNP training, all the hyperparameters were set as previously mentioned, except for a longer radial cutoff of 15 Å and higher maximum $l$ of 2. Finally, we simulated NaCl at 1.3M, 2.6M and 4M using thbe coarse-grained MD simulations with the force field trained on the classical MD simulation of NaCl at 4M. All other hyper parameters for the CG-NNP trained on CMD simulations were the same as the hyperparameters for the CG-NNP trained on all atom NNP-MD.